I am trying to solve the following algebraic problem.
14r/s for r= 32 and s = 8
14r/s = _______?

Answer by stanbon(75887)

solve the following algebraic problem.
14r/s for r= 32 and s = 8
14r/s
= 14*32/8
= 14*4
= 56
=================
Cheers,
Stan H.